# Provable Benefit of Annealed Langevin Monte Carlo for Non-log-concave Sampling

### v1

@@ 6       Accept $X$ as a sample of $\pi_{0}$ when
$U\leq\exp\left(-V_{0}(X)+V_{0}(x^{\prime})+\left\langle\nabla V_{0}(x^{\prime}),X-x^{\prime}\right\rangle+\frac{\lambda_{0}-\eta_{0}\beta}{2}\left\|X-x^{\prime}\right\|^{2}\right).$
@@
Appendix B Proofs for Annealed LMC
B.1 Proof of Equation 7
Define $\varLambda(t):=\int_{0}^{t}\lambda\left(\frac{\tau}{T}\right){\rm d}\tau$, whose derivative is $\varLambda^{\prime}(t)=\lambda\left(\frac{t}{T}\right)$. By Itô’s formula, we have
${\rm d}\left({\rm e}^{\varLambda(t)}X_{t}\right)={\rm e}^{\varLambda(t)}\left(\lambda\left(\frac{t}{T}\right)X_{t}{\rm d}t+{\rm d}X_{t}\right)={\rm e}^{\varLambda(t)}\left(-\eta\left(\frac{t}{T}\right)\nabla V(X_{t_{-}}){\rm d}t+\sqrt{2}{\rm d}B_{t}\right).$
@@ B.1 Proof of Equation 7
${\rm e}^{\varLambda(T_{\ell})}X_{T_{\ell}}-{\rm e}^{\varLambda(T_{\ell-1})}X_{T_{\ell-1}}$ $=-\left(\int_{T_{\ell-1}}^{T_{\ell}}\eta\left(\frac{t}{T}\right){\rm e}^{\varLambda(t)}{\rm d}t\right)\nabla V(X_{T_{\ell-1}})+\sqrt{2}\int_{T_{\ell-1}}^{T_{\ell}}{\rm e}^{\varLambda(t)}{\rm d}B_{t},$
@@
$X_{T_{\ell}}$ $={\rm e}^{-(\varLambda(T_{\ell})-\varLambda(T_{\ell-1}))}X_{T_{\ell-1}}-\left(\int_{T_{\ell-1}}^{T_{\ell}}\eta\left(\frac{t}{T}\right){\rm e}^{-(\varLambda(T_{\ell})-\varLambda(t))}{\rm d}t\right)\nabla V(X_{T_{\ell-1}})$
$+\sqrt{2}\int_{T_{\ell-1}}^{T_{\ell}}{\rm e}^{-(\varLambda(T_{\ell})-\varLambda(t))}{\rm d}B_{t}.$
@@
$\varLambda(T_{\ell})-\varLambda(t)=\int_{t}^{T_{\ell}}\lambda\left(\frac{\tau}{T}\right){\rm d}\tau=T\int_{t/T}^{T_{\ell}/T}\lambda(u){\rm d}u,$
by defining $\Lambda_{0}(\theta^{\prime},\theta)=\exp\left(-T\int_{\theta}^{\theta^{\prime}}\lambda(u){\rm d}u\right)$, we have ${\rm e}^{-(\varLambda(T_{\ell})-\varLambda(T_{\ell-1}))}=\Lambda_{0}(\theta_{\ell},\theta_{\ell-1})$. Similarly, we can show that
$\int_{T_{\ell-1}}^{T_{\ell}}\eta\left(\frac{t}{T}\right){\rm e}^{-(\varLambda(T_{\ell})-\varLambda(t))}{\rm d}t=\int_{T_{\ell-1}}^{T_{\ell}}\eta\left(\frac{t}{T}\right)\Lambda_{0}\left(\frac{T_{\ell}}{T},\frac{t}{T}\right){\rm d}t=T\int_{\theta_{\ell-1}}^{\theta_{\ell}}\eta(u)\Lambda_{0}(\theta_{\ell},u){\rm d}u,$
and $\sqrt{2}\int_{T_{\ell-1}}^{T_{\ell}}{\rm e}^{-(\varLambda(T_{\ell})-\varLambda(t))}{\rm d}B_{t}$ is a zero-mean Gaussian random vector with covariance
$2\int_{T_{\ell-1}}^{T_{\ell}}{\rm e}^{-2(\varLambda(T_{\ell})-\varLambda(t))}{\rm d}t\cdot I=2\int_{T_{\ell-1}}^{T_{\ell}}\Lambda_{0}^{2}\left(\frac{T_{\ell}}{T},\frac{t}{T}\right){\rm d}t\cdot I=2T\int_{\theta_{\ell-1}}^{\theta_{\ell}}\Lambda_{0}^{2}(\theta_{\ell},u){\rm d}u\cdot I.$
B.2 Proof of Theorem 2
We denote the path measure of ALMC (Equation 6) by $\mathbb{Q}$. Then, $\mathbb{Q}_{T}$, the marginal distribution of $X_{T}$, serves as the output distribution $\nu^{\rm ALMC}$. Similar to the methodology in the proof of Theorem 1, we use $\mathbb{P}$ to denote the reference path measure of Equation 5, in which the vector field $(v_{t})_{t\in[0,T]}$ generates the cure of probability distributions $(\widetilde{\pi}_{t})_{t\in[0,T]}$.
Using the data-processing inequality, it suffices to demonstrate that $\operatorname{KL}\left(\mathbb{P}\middle\|\mathbb{Q}\right)\leq\varepsilon^{2}$. By Girsanov theorem (Lemma 1) and triangle inequality, we have
$\operatorname{KL}\left(\mathbb{P}\middle\|\mathbb{Q}\right)$ $=\frac{1}{4}\int_{0}^{T}\operatorname{\mathbb{E}}_{\mathbb{P}}\left[\left\|\eta\left(\frac{t}{T}\right)\left(\nabla V(X_{t})-\nabla V(X_{t_{-}})\right)-v_{t}(X_{t})\right\|^{2}\right]{\rm d}t$
$\lesssim\int_{0}^{T}\operatorname{\mathbb{E}}_{\mathbb{P}}\left[\eta\left(\frac{t}{T}\right)^{2}\left\|\nabla V(X_{t})-\nabla V(X_{t_{-}})\right\|^{2}+\left\|v_{t}(X_{t})\right\|^{2}\right]{\rm d}t$
$\leq\sum_{\ell=1}^{M}\eta\left(\frac{T_{\ell}}{T}\right)^{2}\beta^{2}\int_{T_{\ell-1}}^{T_{\ell}}\operatorname{\mathbb{E}}_{\mathbb{P}}\left[\left\|X_{t}-X_{t_{-}}\right\|^{2}\right]{\rm d}t+\int_{0}^{T}\left\|v_{t}\right\|_{L^{2}(\widetilde{\pi}_{t})}^{2}{\rm d}t.$
The last inequality arises from the smoothness of $V$ and the increasing property of $\eta(\cdot)$. To bound $\operatorname{\mathbb{E}}_{\mathbb{P}}\left[\left\|X_{t}-X_{t_{-}}\right\|^{2}\right]$, note that under $\mathbb{P}$, for $t\in[T_{\ell-1},T_{\ell})$, we have
$X_{t}-X_{t_{-}}=\int_{T_{\ell-1}}^{t}\left(\nabla\log\widetilde{\pi}_{\tau}+v_{\tau}\right)(X_{\tau}){\rm d}\tau+\sqrt{2}(B_{t}-B_{T_{\ell-1}}).$
@@ B.2 Proof of Theorem 2
$\operatorname{\mathbb{E}}_{\mathbb{P}}\left[\left\|X_{t}-X_{t_{-}}\right\|^{2}\right]$ $\lesssim\operatorname{\mathbb{E}}_{\mathbb{P}}\left[\left\|\int_{T_{\ell-1}}^{t}\left(\nabla\log\widetilde{\pi}_{\tau}+v_{\tau}\right)(X_{\tau}){\rm d}\tau\right\|^{2}\right]+\operatorname{\mathbb{E}}\left[\left\|\sqrt{2}(B_{t}-B_{T_{\ell-1}})\right\|^{2}\right]$
$\lesssim(t-{T_{\ell-1}})\int_{T_{\ell-1}}^{t}\operatorname{\mathbb{E}}_{\mathbb{P}}\left[\left\|(\nabla\log\widetilde{\pi}_{\tau}+v_{\tau})(X_{\tau})\right\|^{2}\right]{\rm d}\tau+d(t-{T_{\ell-1}})$
$\lesssim(t-T_{\ell-1})\int_{T_{\ell-1}}^{t}\left(\left\|\nabla\log\widetilde{\pi}_{\tau}\right\|^{2}_{L^{2}(\widetilde{\pi}_{\tau})}+\left\|v_{\tau}\right\|_{L^{2}(\widetilde{\pi}_{\tau})}^{2}\right){\rm d}\tau+d(t-T_{\ell-1})$
$\lesssim h_{\ell}\int_{T_{\ell-1}}^{T_{\ell}}\left(\left\|\nabla\log\widetilde{\pi}_{\tau}\right\|^{2}_{L^{2}(\widetilde{\pi}_{\tau})}+\left\|v_{\tau}\right\|_{L^{2}(\widetilde{\pi}_{\tau})}^{2}\right){\rm d}\tau+dh_{\ell}.$
@@
$\int_{T_{\ell-1}}^{T_{\ell}}\operatorname{\mathbb{E}}\left[\left\|X_{t}-X_{t_{-}}\right\|^{2}\right]{\rm d}t\lesssim h_{\ell}^{2}\int_{T_{\ell-1}}^{T_{\ell}}\left(\left\|\nabla\log\widetilde{\pi}_{t}\right\|^{2}_{L^{2}(\widetilde{\pi}_{t})}+\left\|v_{t}\right\|_{L^{2}(\widetilde{\pi}_{t})}^{2}\right){\rm d}t+dh_{\ell}^{2}.$
Recall that the potential of $\widetilde{\pi}_{t}$ is $\left(\eta\left(\frac{t}{T}\right)\beta+\lambda\left(\frac{t}{T}\right)\right)$-smooth. By Lemma 4 and the monotonicity of $\eta(\cdot)$ and $\lambda(\cdot)$, we have
$\int_{T_{\ell-1}}^{T_{\ell}}\left\|\nabla\log\widetilde{\pi}_{t}\right\|^{2}_{L^{2}(\widetilde{\pi}_{t})}{\rm d}t$ $\leq\int_{T_{\ell-1}}^{T_{\ell}}d\left(\eta\left(\frac{t}{T}\right)\beta+\lambda\left(\frac{t}{T}\right)\right){\rm d}t$
@@
Therefore, $\operatorname{KL}\left(\mathbb{P}\middle\|\mathbb{Q}\right)$ is upper bounded by
$\lesssim\sum_{\ell=1}^{M}\eta\left(\theta_{\ell}\right)^{2}\beta^{2}\int_{T_{\ell-1}}^{T_{\ell}}\operatorname{\mathbb{E}}_{\mathbb{P}}\left[\left\|X_{t}-X_{t_{-}}\right\|^{2}\right]{\rm d}t+\int_{0}^{T}\left\|v_{t}\right\|_{L^{2}(\widetilde{\pi}_{t})}^{2}{\rm d}t$
$\lesssim\sum_{\ell=1}^{M}\eta\left(\theta_{\ell}\right)^{2}\beta^{2}\left(dh_{\ell}^{3}\left(\beta\eta\left(\theta_{\ell}\right)+\lambda\left(\theta_{\ell-1}\right)\right)+h_{\ell}^{2}\int_{T_{\ell-1}}^{T_{\ell}}\left\|v_{t}\right\|_{L^{2}(\widetilde{\pi}_{t})}^{2}{\rm d}t+dh_{\ell}^{2}\right)+\int_{0}^{T}\left\|v_{t}\right\|_{L^{2}(\widetilde{\pi}_{t})}^{2}{\rm d}t$
$=\sum_{\ell=1}^{M}\left(\left(1+\eta(\theta_{\ell})^{2}\beta^{2}h_{\ell}^{2}\right)\int_{T_{\ell-1}}^{T_{\ell}}\left\|v_{t}\right\|_{L^{2}(\widetilde{\pi}_{t})}^{2}{\rm d}t+\eta(\theta_{\ell})^{2}\beta^{2}dh_{\ell}^{2}\left(1+h_{\ell}\left(\beta\eta(\theta_{\ell})+\lambda(\theta_{\ell-1})\right)\right)\right).$
For the remaining integral, given that $(v_{t})_{t\in[0,T]}$ generates $(\widetilde{\pi}_{t})_{t\in[0,T]}$, according to Lemma 2, we may choose $v_{t}$ such that $\left\|v_{t}\right\|_{L^{2}(\widetilde{\pi}_{t})}=\left|\dot{\widetilde{\pi}}\right|_{t}$. Thus,
$\int_{T_{\ell-1}}^{T_{\ell}}\left\|v_{t}\right\|_{L^{2}(\widetilde{\pi}_{t})}^{2}{\rm d}t$ $=\int_{T_{\ell-1}}^{T_{\ell}}\left|\dot{\widetilde{\pi}}\right|_{t}^{2}{\rm d}t=\int_{T_{\ell-1}}^{T_{\ell}}\frac{1}{T^{2}}\left|\dot{\pi}\right|^{2}_{t/T}{\rm d}t=\frac{1}{T}\int_{\theta_{\ell-1}}^{\theta_{\ell}}\left|\dot{\pi}\right|^{2}_{\theta}{\rm d}\theta,$
through a change-of-variable analogous to that used in the proof of Theorem 1. Therefore,
$\operatorname{KL}\left(\mathbb{P}\middle\|\mathbb{Q}\right)$ $\lesssim\sum_{\ell=1}^{M}\left(\frac{1+\eta(\theta_{\ell})^{2}\beta^{2}h_{\ell}^{2}}{T}\int_{\theta_{\ell-1}}^{\theta_{\ell}}\left|\dot{\pi}\right|^{2}_{\theta}{\rm d}\theta+\eta(\theta_{\ell})^{2}\beta^{2}dh_{\ell}^{2}\left(1+h_{\ell}\left(\beta\eta(\theta_{\ell})+\lambda(\theta_{\ell-1})\right)\right)\right).$
@@
$\operatorname{KL}\left(\mathbb{P}\middle\|\mathbb{Q}\right)$ $\lesssim\sum_{\ell=1}^{M}\left(\frac{1}{T}\int_{\theta_{\ell-1}}^{\theta_{\ell}}\left|\dot{\pi}\right|^{2}_{\theta}{\rm d}\theta+\eta(\theta_{\ell})^{2}\beta^{2}dh_{\ell}^{2}\right)$
$=\frac{{\cal A}_{\eta,\lambda}(\pi)}{T}+\beta^{2}d\sum_{\ell=1}^{M}\eta(\theta_{\ell})^{2}h_{\ell}^{2}$
$=\frac{{\cal A}_{\eta,\lambda}(\pi)}{T}+\beta^{2}d\sum_{\ell=1}^{M}\eta(\theta_{\ell})^{2}T^{2}(\theta_{\ell}-\theta_{\ell-1})^{2}.$
To bound the above expression by $\varepsilon^{2}$, we first select $T\asymp\frac{{\cal A}_{\eta,\lambda}(\pi)}{\varepsilon^{2}}$, mirroring the total time $T$ required for running annealed LD as specified in Theorem 1. This guarantees that the continuous dynamics closely approximate the reference path measure. Given that $\eta(\cdot)\leq 1$, it remains only to ensure
$\beta^{2}d\sum_{\ell=1}^{M}\eta(\theta_{\ell})^{2}T^{2}(\theta_{\ell}-\theta_{\ell-1})^{2}\leq\beta^{2}d\frac{{\cal A}_{\eta,\lambda}(\pi)^{2}}{\varepsilon^{4}}\sum_{\ell=1}^{M}(\theta_{\ell}-\theta_{\ell-1})^{2}\lesssim\varepsilon^{2},$
@@
$\sum_{\ell=1}^{M}(\theta_{\ell}-\theta_{\ell-1})^{2}\lesssim\frac{\varepsilon^{6}}{d\beta^{2}{\cal A}_{\eta,\lambda}(\pi)^{2}}.$ (8)
@@
$M\asymp\frac{d\beta^{2}{\cal A}_{\eta,\lambda}(\pi)^{2}}{\varepsilon^{6}}$
satisfies the constraint given in Equation 8. In this case, the step size $h_{\ell}=\frac{T}{M}\asymp\frac{\varepsilon^{4}}{d\beta^{2}}\lesssim\frac{1}{\beta d}$. Combining this with the $\widetilde{O}\left(1\right)$ complexity for sampling from $\pi_{0}$, we have completed the proof.
Appendix C Proof of Example 1
The smoothness of $V$ comes from [cheng2023fast, Lemma 4].
Note that $\pi(x)\propto\sum_{i=1}^{N}p_{i}\exp\left(-\frac{\beta}{2}\left\|x-y_{i}\right\|^{2}\right)$, and define
$\widehat{\pi}_{\lambda}(x)$ $\propto\pi(x)\exp\left(-\frac{\lambda}{2}\left\|x\right\|^{2}\right)$
$=\sum_{i=1}^{N}p_{i}\exp\left(-\frac{\lambda\beta}{2(\lambda+\beta)}\left\|y_{i}\right\|^{2}-\frac{\lambda+\beta}{2}\left\|x-\frac{\beta}{\lambda+\beta}y_{i}\right\|^{2}\right)$
@@ Appendix C Proof of Example 1
We use coupling method to upper bound $W_{2}^{2}(\widehat{\pi}_{\lambda},\widehat{\pi}_{\lambda+\delta})$. We first sample $I$ with distribution $\operatorname{\mathbb{P}}\left(I=i\right)=p_{i}$, $i\in\left[\!\left[1,N\right]\!\right]$, and then independently sample $\eta\sim\operatorname{{\cal N}}\left(0,I\right)$. Then,

### v2

@@ 6       Accept $X$ as a sample of $\pi_{0}$ when
$U\leq\exp\left(-V_{0}(X)+V_{0}(x^{\prime})+\left\langle\nabla V_{0}(x^{\prime}),X-x^{\prime}\right\rangle+\frac{\lambda_{0}-\eta_{0}\beta}{2}\|x-x^{\prime}\|^{2}\right).$
@@
Remark.
The parameter $R$ reflects prior knowledge about global minimizer(s) of the potential function $V$. Unless it is exceptionally large, indicating scarce prior information about the global minimizer(s) of $V$, this $\widetilde{O}(1)$ complexity is negligible compared to the overall complexity of sampling. In particular, when the exact location of a global minimizer of $V$ is known, we can adjust the potential $V$ so that 00 becomes a global minimizer, thereby reducing the complexity to $O(1)$.
Appendix C Proofs for Annealed LMC
C.1 Proof of Eq. 7
Define $\varLambda(t):=\int_{0}^{t}\lambda\left(\frac{\tau}{T}\right)\mathrm{d}\tau$, whose derivative is $\varLambda^{\prime}(t)=\lambda\left(\frac{t}{T}\right)$. By Itô’s formula, we have
$\mathrm{d}\left({\rm e}^{\varLambda(t)}X_{t}\right)={\rm e}^{\varLambda(t)}\left(\lambda\left(\frac{t}{T}\right)X_{t}\mathrm{d}t+\mathrm{d}X_{t}\right)={\rm e}^{\varLambda(t)}\left(-\eta\left(\frac{t}{T}\right)\nabla V(X_{t_{-}})\mathrm{d}t+\sqrt{2}\mathrm{d}B_{t}\right).$
@@ Appendix C Proofs for Annealed LMC
${\rm e}^{\varLambda(T_{\ell})}X_{T_{\ell}}-{\rm e}^{\varLambda(T_{\ell-1})}X_{T_{\ell-1}}$ $=-\left(\int_{T_{\ell-1}}^{T_{\ell}}\eta\left(\frac{t}{T}\right){\rm e}^{\varLambda(t)}\mathrm{d}t\right)\nabla V(X_{T_{\ell-1}})+\sqrt{2}\int_{T_{\ell-1}}^{T_{\ell}}{\rm e}^{\varLambda(t)}\mathrm{d}B_{t},$
@@
$X_{T_{\ell}}$ $={\rm e}^{-(\varLambda(T_{\ell})-\varLambda(T_{\ell-1}))}X_{T_{\ell-1}}-\left(\int_{T_{\ell-1}}^{T_{\ell}}\eta\left(\frac{t}{T}\right){\rm e}^{-(\varLambda(T_{\ell})-\varLambda(t))}\mathrm{d}t\right)\nabla V(X_{T_{\ell-1}})$
$+\sqrt{2}\int_{T_{\ell-1}}^{T_{\ell}}{\rm e}^{-(\varLambda(T_{\ell})-\varLambda(t))}\mathrm{d}B_{t}.$
@@
$\varLambda(T_{\ell})-\varLambda(t)=\int_{t}^{T_{\ell}}\lambda\left(\frac{\tau}{T}\right)\mathrm{d}\tau=T\int_{t/T}^{T_{\ell}/T}\lambda(u)\mathrm{d}u,$
by defining $\Lambda_{0}(\theta^{\prime},\theta)=\exp\left(-T\int_{\theta}^{\theta^{\prime}}\lambda(u)\mathrm{d}u\right)$, we have ${\rm e}^{-(\varLambda(T_{\ell})-\varLambda(T_{\ell-1}))}=\Lambda_{0}(\theta_{\ell},\theta_{\ell-1})$. Similarly, we can show that
$\int_{T_{\ell-1}}^{T_{\ell}}\eta\left(\frac{t}{T}\right){\rm e}^{-(\varLambda(T_{\ell})-\varLambda(t))}\mathrm{d}t=\int_{T_{\ell-1}}^{T_{\ell}}\eta\left(\frac{t}{T}\right)\Lambda_{0}\left(\frac{T_{\ell}}{T},\frac{t}{T}\right)\mathrm{d}t=T\int_{\theta_{\ell-1}}^{\theta_{\ell}}\eta(u)\Lambda_{0}(\theta_{\ell},u)\mathrm{d}u,$
and $\sqrt{2}\int_{T_{\ell-1}}^{T_{\ell}}{\rm e}^{-(\varLambda(T_{\ell})-\varLambda(t))}\mathrm{d}B_{t}$ is a zero-mean Gaussian random vector with covariance
$2\int_{T_{\ell-1}}^{T_{\ell}}{\rm e}^{-2(\varLambda(T_{\ell})-\varLambda(t))}\mathrm{d}t\cdot I=2\int_{T_{\ell-1}}^{T_{\ell}}\Lambda_{0}^{2}\left(\frac{T_{\ell}}{T},\frac{t}{T}\right)\mathrm{d}t\cdot I=2T\int_{\theta_{\ell-1}}^{\theta_{\ell}}\Lambda_{0}^{2}(\theta_{\ell},u)\mathrm{d}u\cdot I.$
$\square$
C.2 Proof of Thm. 2
We denote the path measure of ALMC (Eq. 6) by $\mathbb{Q}$. Then, $\mathbb{Q}_{T}$, the marginal distribution of $X_{T}$, serves as the output distribution $\nu^{\rm ALMC}$. Similar to the methodology in the proof of Thm. 1, we use $\mathbb{P}$ to denote the reference path measure of Eq. 4, in which the vector field $(v_{t})_{t\in[0,T]}$ generates the cure of probability distributions $(\widetilde{\pi}_{t})_{t\in[0,T]}$.
Using the data-processing inequality, it suffices to demonstrate that $\operatorname{KL}(\mathbb{P}\|\mathbb{Q})\leq\varepsilon^{2}$. By Girsanov theorem (Lem. 1) and triangle inequality, we have
$\operatorname{KL}(\mathbb{P}\|\mathbb{Q})$ $=\frac{1}{4}\int_{0}^{T}\mathbb{E}_{\mathbb{P}}{\left\|\eta\left(\frac{t}{T}\right)\left(\nabla V(X_{t})-\nabla V(X_{t_{-}})\right)-v_{t}(X_{t})\right\|^{2}}\mathrm{d}t$
$\lesssim\int_{0}^{T}\mathbb{E}_{\mathbb{P}}{\left[\eta\left(\frac{t}{T}\right)^{2}\|\nabla V(X_{t})-\nabla V(X_{t_{-}})\|^{2}+\|v_{t}(X_{t})\|^{2}\right]}\mathrm{d}t$
$\leq\sum_{\ell=1}^{M}\eta\left(\frac{T_{\ell}}{T}\right)^{2}\beta^{2}\int_{T_{\ell-1}}^{T_{\ell}}\mathbb{E}_{\mathbb{P}}{\|X_{t}-X_{t_{-}}\|^{2}}\mathrm{d}t+\int_{0}^{T}\|v_{t}\|_{L^{2}(\widetilde{\pi}_{t})}^{2}\mathrm{d}t.$
The last inequality arises from the smoothness of $V$ and the increasing property of $\eta(\cdot)$. To bound $\mathbb{E}_{\mathbb{P}}{\|X_{t}-X_{t_{-}}\|^{2}}$, note that under $\mathbb{P}$, for $t\in[T_{\ell-1},T_{\ell})$, we have
$X_{t}-X_{t_{-}}=\int_{T_{\ell-1}}^{t}\left(\nabla\log\widetilde{\pi}_{\tau}+v_{\tau}\right)(X_{\tau})\mathrm{d}\tau+\sqrt{2}(B_{t}-B_{T_{\ell-1}}).$
@@
$\mathbb{E}_{\mathbb{P}}{\|X_{t}-X_{t_{-}}\|^{2}}$ $\lesssim\mathbb{E}_{\mathbb{P}}{\left\|\int_{T_{\ell-1}}^{t}\left(\nabla\log\widetilde{\pi}_{\tau}+v_{\tau}\right)(X_{\tau})\mathrm{d}\tau\right\|^{2}}+\mathbb{E}{\|\sqrt{2}(B_{t}-B_{T_{\ell-1}})\|^{2}}$
$\lesssim(t-{T_{\ell-1}})\int_{T_{\ell-1}}^{t}\mathbb{E}_{\mathbb{P}}{\|(\nabla\log\widetilde{\pi}_{\tau}+v_{\tau})(X_{\tau})\|^{2}}\mathrm{d}\tau+d(t-{T_{\ell-1}})$
$\lesssim(t-T_{\ell-1})\int_{T_{\ell-1}}^{t}\left(\|\nabla\log\widetilde{\pi}_{\tau}\|^{2}_{L^{2}(\widetilde{\pi}_{\tau})}+\|v_{\tau}\|_{L^{2}(\widetilde{\pi}_{\tau})}^{2}\right)\mathrm{d}\tau+d(t-T_{\ell-1})$
$\lesssim h_{\ell}\int_{T_{\ell-1}}^{T_{\ell}}\left(\|\nabla\log\widetilde{\pi}_{\tau}\|^{2}_{L^{2}(\widetilde{\pi}_{\tau})}+\|v_{\tau}\|_{L^{2}(\widetilde{\pi}_{\tau})}^{2}\right)\mathrm{d}\tau+dh_{\ell}.$
@@
$\int_{T_{\ell-1}}^{T_{\ell}}\mathbb{E}{\|X_{t}-X_{t_{-}}\|^{2}}\mathrm{d}t\lesssim h_{\ell}^{2}\int_{T_{\ell-1}}^{T_{\ell}}\left(\|\nabla\log\widetilde{\pi}_{t}\|^{2}_{L^{2}(\widetilde{\pi}_{t})}+\|v_{t}\|_{L^{2}(\widetilde{\pi}_{t})}^{2}\right)\mathrm{d}t+dh_{\ell}^{2}.$
Recall that the potential of $\widetilde{\pi}_{t}$ is $\left(\eta\left(\frac{t}{T}\right)\beta+\lambda\left(\frac{t}{T}\right)\right)$-smooth. By Lem. 6 and the monotonicity of $\eta(\cdot)$ and $\lambda(\cdot)$, we have
$\int_{T_{\ell-1}}^{T_{\ell}}\|\nabla\log\widetilde{\pi}_{t}\|^{2}_{L^{2}(\widetilde{\pi}_{t})}\mathrm{d}t$ $\leq\int_{T_{\ell-1}}^{T_{\ell}}d\left(\eta\left(\frac{t}{T}\right)\beta+\lambda\left(\frac{t}{T}\right)\right)\mathrm{d}t$
@@
Therefore, $\operatorname{KL}(\mathbb{P}\|\mathbb{Q})$ is upper bounded by
$\lesssim\sum_{\ell=1}^{M}\eta\left(\theta_{\ell}\right)^{2}\beta^{2}\int_{T_{\ell-1}}^{T_{\ell}}\mathbb{E}_{\mathbb{P}}{\|X_{t}-X_{t_{-}}\|^{2}}\mathrm{d}t+\int_{0}^{T}\|v_{t}\|_{L^{2}(\widetilde{\pi}_{t})}^{2}\mathrm{d}t$
$\lesssim\sum_{\ell=1}^{M}\eta\left(\theta_{\ell}\right)^{2}\beta^{2}\left(dh_{\ell}^{3}\left(\beta\eta\left(\theta_{\ell}\right)+\lambda\left(\theta_{\ell-1}\right)\right)+h_{\ell}^{2}\int_{T_{\ell-1}}^{T_{\ell}}\|v_{t}\|_{L^{2}(\widetilde{\pi}_{t})}^{2}\mathrm{d}t+dh_{\ell}^{2}\right)+\int_{0}^{T}\|v_{t}\|_{L^{2}(\widetilde{\pi}_{t})}^{2}\mathrm{d}t$
$=\sum_{\ell=1}^{M}\left(\left(1+\eta(\theta_{\ell})^{2}\beta^{2}h_{\ell}^{2}\right)\int_{T_{\ell-1}}^{T_{\ell}}\|v_{t}\|_{L^{2}(\widetilde{\pi}_{t})}^{2}\mathrm{d}t+\eta(\theta_{\ell})^{2}\beta^{2}dh_{\ell}^{2}\left(1+h_{\ell}\left(\beta\eta(\theta_{\ell})+\lambda(\theta_{\ell-1})\right)\right)\right).$
For the remaining integral, given that $(v_{t})_{t\in[0,T]}$ generates $(\widetilde{\pi}_{t})_{t\in[0,T]}$, according to Lem. 2, we may choose $v_{t}$ such that $\|v_{t}\|_{L^{2}(\widetilde{\pi}_{t})}=|\dot{\widetilde{\pi}}|_{t}$. Thus,
$\int_{T_{\ell-1}}^{T_{\ell}}\|v_{t}\|_{L^{2}(\widetilde{\pi}_{t})}^{2}\mathrm{d}t$ $=\int_{T_{\ell-1}}^{T_{\ell}}|\dot{\widetilde{\pi}}|_{t}^{2}\mathrm{d}t=\int_{T_{\ell-1}}^{T_{\ell}}\frac{1}{T^{2}}|\dot{\pi}|^{2}_{t/T}\mathrm{d}t=\frac{1}{T}\int_{\theta_{\ell-1}}^{\theta_{\ell}}|\dot{\pi}|^{2}_{\theta}\mathrm{d}\theta,$
through a change-of-variable analogous to that used in the proof of Thm. 1. Therefore,
$\operatorname{KL}(\mathbb{P}\|\mathbb{Q})$ $\lesssim\sum_{\ell=1}^{M}\left(\frac{1+\eta(\theta_{\ell})^{2}\beta^{2}h_{\ell}^{2}}{T}\int_{\theta_{\ell-1}}^{\theta_{\ell}}|\dot{\pi}|^{2}_{\theta}\mathrm{d}\theta+\eta(\theta_{\ell})^{2}\beta^{2}dh_{\ell}^{2}\left(1+h_{\ell}\left(\beta\eta(\theta_{\ell})+\lambda(\theta_{\ell-1})\right)\right)\right).$
@@
$\operatorname{KL}(\mathbb{P}\|\mathbb{Q})$ $\lesssim\sum_{\ell=1}^{M}\left(\frac{1}{T}\int_{\theta_{\ell-1}}^{\theta_{\ell}}|\dot{\pi}|^{2}_{\theta}\mathrm{d}\theta+\eta(\theta_{\ell})^{2}\beta^{2}dh_{\ell}^{2}\right)$
$=\frac{{\cal A}}{T}+\beta^{2}d\sum_{\ell=1}^{M}\eta(\theta_{\ell})^{2}h_{\ell}^{2}$
$=\frac{{\cal A}}{T}+\beta^{2}d\sum_{\ell=1}^{M}\eta(\theta_{\ell})^{2}T^{2}(\theta_{\ell}-\theta_{\ell-1})^{2}.$
To bound the above expression by $\varepsilon^{2}$, we first select $T\asymp\frac{{\cal A}}{\varepsilon^{2}}$, mirroring the total time $T$ required for running annealed LD as specified in Thm. 1. This guarantees that the continuous dynamics closely approximate the reference path measure. Given that $\eta(\cdot)\leq 1$, it remains only to ensure
$\beta^{2}d\sum_{\ell=1}^{M}\eta(\theta_{\ell})^{2}T^{2}(\theta_{\ell}-\theta_{\ell-1})^{2}\leq\beta^{2}d\frac{{\cal A}^{2}}{\varepsilon^{4}}\sum_{\ell=1}^{M}(\theta_{\ell}-\theta_{\ell-1})^{2}\lesssim\varepsilon^{2},$
@@
$\sum_{\ell=1}^{M}(\theta_{\ell}-\theta_{\ell-1})^{2}\lesssim\frac{\varepsilon^{6}}{d\beta^{2}{\cal A}^{2}}.$ (8)
@@
$M\asymp\frac{d\beta^{2}{\cal A}^{2}}{\varepsilon^{6}}$
satisfies the constraint given in Eq. 8. In this case, the step size $h_{\ell}=\frac{T}{M}\asymp\frac{\varepsilon^{4}}{d\beta^{2}}\lesssim\frac{1}{\beta d}$. Combining this with the $\widetilde{O}(1)$ complexity for sampling from $\pi_{0}$, we have completed the proof. $\square$
Appendix D Proof of Ex. 2
The smoothness of $V$ comes from Lem. 7.
Note that $\pi(x)\propto\sum_{i=1}^{N}p_{i}\exp\left(-\frac{\beta}{2}\|x-y_{i}\|^{2}\right)$, and define
$\widehat{\pi}_{\lambda}(x)$ $\propto\pi(x)\exp\left(-\frac{\lambda}{2}\|x\|^{2}\right)$
$=\sum_{i=1}^{N}p_{i}\exp\left(-\frac{\lambda\beta}{2(\lambda+\beta)}\|y_{i}\|^{2}-\frac{\lambda+\beta}{2}\left\|x-\frac{\beta}{\lambda+\beta}y_{i}\right\|^{2}\right)$
@@
We use coupling method to upper bound $W_{2}^{2}(\widehat{\pi}_{\lambda},\widehat{\pi}_{\lambda+\delta})$. We first sample $I$ following distribution $\mathbb{P}(I=i)=p_{i}$, $i\in\left[\!\left[1,N\right]\!\right]$, and then independently sample $\eta\sim\operatorname{{\cal N}}\left(0,I\right)$. We have